• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name E.Chudakov

    Log entry time 23:57:06 on January11,2002

    Entry number 80114

    keyword=Shift summary

    Beam tune started at about 20:00
    High radiation detected on the target ion chamber, initially about 1krad/h at 1uA CW
    Various equipment elements tripped, all but ROC4 have been reset. Fixing ROC4 needs an access to the hall.
    MCC continues working on beam tunning.